# Recidivism and Peer Influence with LLM Text Embeddings in Low Security Correctional Facilities

### v1

@@ 3.6 Network models
In terms of latent variable network models, we first consider the Latent Space Model (LSM), which is a general random graph model that includes both additive and multiplicative latent variables and can also accommodate covariates (Hoff et al., 2002; Hoff, 2021; Ma et al., 2020; Li et al., 2023). The model is parameterized by $d$-dimensional unknown vectors $\bm{q}_{i}$ and unknown scalars $v_{i}$ for all nodes $i=1,\ldots,n$.
We also assume that we have edge level covariates $\bm{X}=(x_{ij})$ available to us. A network adjacency matrix $\bm{A}=(a_{ij})$ from this model is generated as follows (Li et al., 2023; Ma et al., 2020; Chang and Paul, 2024):
@@
where $f$ is a family of distributions satisfying certain smoothness conditions, and $\sigma:\mathbb{R}\to\mathbb{R}$ is a known link function. In this paper, we consider a particular sparse version of the model due to Li et al. (2023) that assumes $f$ to be a Bernoulli distribution, link function $\sigma$ to be the logistic function, and there exists a sparsity parameter $\rho_{N}$ such that $\theta_{ij}=logistic({\bm{q}_{i}}^{\prime}\bm{q}_{j}+v_{i}+v_{j}+x_{ij}\beta+\rho_{N})$. Defining $\omega_{N}=\exp(\rho_{N})$, we assume $\omega_{N}\to 0$ and $\omega_{N}=\omega(N^{-1/2})$. Let $\bm{u}_{i}\triangleq(\bm{q}_{i}^{\prime},v_{i})^{\prime}$ denote a parameter vector containing all latent variables associated with node $i$. Then, we can write the latent variable matrix collectively as
$\bm{U}:=[\bm{u}_{1},\dots,\bm{u}_{N}]^{\prime}=\bm{Q}\bm{Q}^{\prime}+\bm{v}\bm{1}_{N}^{\prime}+\bm{1}_{N}\bm{v}^{\prime}.$
A special case of this model is Random Dot Product Graph (RDPG) model, which can be expressed as follows,
@@
where $\rho_{N}$ again controls the sparsity of the model. We will assume $\rho_{N}=\omega(\frac{\log^{4}N}{N})$. Note the expected density (i.e., probability of an edge) scales with $N$ as $\rho_{N}$, or in other words, the expected degree scales as $N\rho_{N}$ in this model, making the model suitable for sparse networks. This model does not account for the additive latent variables $v_{i}$’s but contains $d$-dimensional multiplicative latent variables, and the link function $\sigma(\cdot)$ is specified as identity. Clearly, both these models allow for multidimensional latent variables in the network formation model and is more general than the single latent variable model considered in Johnsson and Moon (2021). Both models are also capable of modeling sparse networks, which the models in Johnsson and Moon (2021); Graham (2017); Auerbach (2022) are not capable of. We compare the model fit from these network models with the network model considered in Johnsson and Moon (2021); Graham (2017) for the network data in our application in real data analysis (see section 5.2).
We employ the maximum likelihood estimator with Lagrange adjustment in Li et al. (2023) for the estimation of the LSM model. Let $\hat{\bm{U}}$ denote the maximum likelihood estimator over the constrained parameter space
@@
Under the set of assumptions laid out in Li et al. (2023), an upper bound on the error rate of estimating the latent positions in Frobenius norm (Li et al., 2023) is $\frac{1}{N}\|\hat{\bm{U}}-\bm{U}\|_{F}^{2}=O_{p}(\frac{1}{N\omega_{N}})$.
Although the LSM described above contains the RDPG model as its special case, there exists a rich amount of results with better convergence rates of the estimate of $\bm{U}$ when one uses spectral embedding to approximate $\bm{U}$ in the case of RDPG (Athreya et al., 2017; Cape et al., 2019; Xie and Xu, 2023). Specifically, the spectral embedding of the adjacency matrix $A$ can be written as $\hat{\bm{U}}=\bm{U}_{A}|\bm{S}_{A}|^{1/2}$, where $\bm{U}_{A}$ denotes the matrix containing the leading $d$ eigenvectors of $\bm{A}$, which are assumed, without loss of generality, to be ordered by decreasing absolute eigenvalue magnitude. The matrix $|\bm{S}_{A}|$ is a diagonal matrix containing the absolute values of the corresponding eigenvalues. Then, a faster convergence rate than LSM can be obtained in terms of the $2\to\infty$ norm as (Rubin-Delanchy et al., 2022) $\|\hat{\bm{U}}-\bm{U}\bm{H}\|_{2,\infty}=O_{hp}\left(\frac{\log^{c}N}{N^{1/2}}\right)$. Here we mean $X_{n}=O_{hp}(1)$ by: for every $c>0$ there exist constants $M(c),n_{0}(c)>0$ such that $\Pr(|X_{n}|>M)<n^{-c}$ for all $n>n_{0}$.
@@ 3.7 Theoretical results
For a matrix $A$ we denote its spectral norm as $\|A\|=\max_{\|x\|=1}|Ax|$, Frobenius norm as $\|A\|_{F}=\sqrt{\sum_{ij}a_{ij}^{2}}$ and the $2\to\infty$ norm as $\|A\|_{2,\infty}=\max_{i}\sqrt{\sum_{j}a_{ij}^{2}}$. We use the stochastic order notation $o_{p}(1)$ to mean that if $x_{N}=o_{p}(1)$, then $x_{N}\overset{p}{\to}0$. We start with a list of assumptions. The first assumption is a collection of conditions described above in the description of the model, making the model well-behaved and identification possible, similar to those in (Johnsson and Moon, 2021).
@@
We note again that the network $\bm{A}$ is endogenous by the dependence between $\bm{U}_{i}$ and $\bm{E}_{i}$. In addition, this implies that
@@
The above assumption is an extension of the assumption in Johnsson and Moon (2021) to multivariate basis functions which says the true component-wise functions $h_{j}(u)$ are well approximated by the sieve approximations for all vectors $u$ and all components $j$.
@@
Assume the assumptions in the section 3.7. Let the graph $\bm{A}$ be generated via either RDPG or LSM as described in section 3.6, respectively. Further assume that
@@
The proofs of all results are in the Appendix A. Several additional technical lemmas are needed, which are all stated and proved in the Appendix A.
Algorithm 1 Vectorized 2SLS with Network Endogeneity in MSAR
[…]
1:Step 1: Spectral Embedding (RDPG)
[…]
3:Step 2: Construct Peer Lag Matrices
[…]
6:Step 3: Latent Adjustment
[…]
10:Step 4: Vectorization
[…]
14:Step 5: IV Estimation and Covariance Calculation
[…]
Note that the conditions on the graph density indicate that consistent and asymptotically normal estimation is possible even for sparse graphs. Under a binary graph ($a_{ij}$s are either 1 or 0), for example, the first condition in (6) only requires $\delta_{\max}/\delta_{\min}^{2}=o_{p}(1/\sqrt{N})$. A sufficient condition for the second part in (7) can be
@@
These conditions hold under a moderate expected density of the graph. For example, if we let $\delta_{\min}\asymp\delta_{\max}$ and the average expected density of the graph grows as $O(\frac{1}{\sqrt{N}})$, then all conditions are satisfied. This is in contrast to the results in Johnsson and Moon (2021), which only hold if the average expected density of the graph grows as $O(1)$, making the graph unrealistically dense.
@@
Let Assumption 3 hold. Under the conditions on RDPG model in Rubin-Delanchy et al. (2022) and sparsity parameter $\rho_{N}=\omega(\frac{\log^{4}N}{N})$, we have
@@
and under the conditions on the sparse LSM model with sparsity parameter $\omega_{N}=\omega(N^{-1/2})$ in Li et al. (2023), we have
@@
Table 1: Monte Carlo Simulations RDPG Model
Y1 Equation Y2 Equation
N $D_{11}$ $D_{12}$ $D_{21}$ $D_{22}$
$h^{Y}(U)=U\Theta_{Y}$
100 1.349 0.284 1.662 0.383
300 1.057 0.224 1.041 0.217
500 0.471 0.098 0.472 0.100
$h^{Y}(U)=[U[,1],U[,2],U[,1]^{2},U[,2]^{2}]\Theta_{Y}$
100 0.638 0.218 0.777 0.279
300 0.256 0.090 0.157 0.055
500 0.205 0.073 0.113 0.039
$h^{Y}(U)=[U[,1],U[,2],U[,1]^{2},U[,2]^{2},U[,1]*U[,2]]\Theta_{Y}$
100 1.671 0.635 1.167 0.460
300 0.514 0.195 0.693 0.258
500 0.156 0.062 0.220 0.091
@@ 4 Simulations
Table 1 documents the performance of the estimator when the network is generated from the RDPG model and the spectral embedding is used to estimate the latent variables. The dimension of latent variables in the outcome model is fixed at 2. There are two correlated outcomes where the data generation follows Equation 1 with endogeneity in network formation. The peer influence parameter matrix is set at $\begin{bmatrix}0.3&0.2\\ 0.25&0.6\end{bmatrix}$ where the diagonal elements capture the direct peer influence spillovers and the off-diagonal elements capture the indirect spillovers in the outcome. The dimension of the coefficients on $h^{Y}(U)$ varies across the three scenarios. For each scenario, the data is generated 100 times, followed by estimation of $\beta_{v}$ using the steps in algorithm 1. The table shows that the mean squared error steadily declines with the number of nodes in the network. This supports the fact that the estimator works well in recovering the true parameter of interest as the sample size increases.
Table 2: Monte Carlo Simulations Latent Space Model with Covariates
Y1 Equation Y2 Equation
N $D_{11}$ $D_{12}$ $D_{21}$ $D_{22}$
$h^{Y}(U)=U\Theta_{Y}$
100 2.434 3.711 1.489 2.241
300 0.874 1.169 0.686 1.064
500 0.413 0.644 0.666 0.833
$h^{Y}(U)=[U[,1],U[,2],U[,1]^{2},U[,2]^{2}]\Theta_{Y}$
100 1.362 1.700 4.380 5.600
300 0.670 0.861 3.957 5.093
500 0.263 0.338 0.987 1.268
$h^{Y}(U)=[U[,1],U[,2],U[,1]^{2},U[,2]^{2},U[,1]*U[,2]]\Theta_{Y}$
100 1.820 2.261 3.256 4.055
300 1.120 1.423 1.933 2.478
500 0.605 0.769 0.838 1.072
Next, we change the data generation and estimation of latent variables, where latent space models that allow for both multiplicative and additive unobserved variables, along with observed covariates, are used for generating the network. A Projected Gradient Descent (PGD) algorithm with Universal Singular Value Thresholding (USVT) initialization is used to estimate the MLE $\hat{U}$ (Ma et al., 2020). Note that even though this solves a non-convex optimization problem, the solution has nice theoretical guarantees which we have used in the proofs above. The covariate that explains the network formation is not part of the data generation for the outcomes. The network formation consists of one additive latent factor, two multiplicative latent factors, and two observed covariates. The outcome data generation consists of $AY,X,AX,h^{Y}(U)]$ where $AY$ is $N\times 2$ matrix, $X$ is $N\times 5$, $AX$ is $N\times 5$ and dimension of $h^{Y}(U)$ varies across the 3 scenarios. The peer influence matrix $D$ used in the data generation for the latent space model with covariates is fixed at $\begin{bmatrix}0.8&0.2\\ 0.3&0.6\end{bmatrix}$. Table 2 reports the mean squared error (MSE) across 50 simulations for each scenario as the sample size increases from 100 to 300 to 500. It shows that the MSE consistently declines as the sample size increases for all four parameters of the peer influence $D$ matrix for all three scenarios.
@@ 5.1 Predict Recidivism using Text Messages
Figure 5 provides the Receiver Operator Characteristic (ROC) curves and the Area Under the Curve (AUC) for two models, one with both pre-entry covariates ($X$) and embeddings ($T$), and an alternative model without the text embeddings (only $X$) as predictors. Panels (a) and (b) show that the prediction accuracy is 30% higher for recidivism in the female unit when we incorporate the LLM-based text embeddings relative to the model with only pre-entry covariates $X$. The dimension reduction is substantial for the former model, as about 26-37 predictors out of text embeddings (768) and pre-entry covariates remain across the five folds in LASSO model fit. Panels (c) and (d) compare the prediction accuracy for the male units with and without the embeddings. The prediction accuracy improvement for the males is 30.6%. Figure B6 repeats the same analysis but for the case when the text embeddings are aggregated by the ID of the receiver rather than the sender in Figure 5. We find comparable improvements (32.6% for females and 32% for males) in predictive accuracy when the text embeddings are aggregated by the receiver profiles and included as predictors in the LASSO model along with the pre-entry covariates. Recall from section 2 that the sender profile indicates how an individual interacts with their peers, while the receiver profiles indicate how others perceive an individual.
@@
a. Female – Without Embeddings
b. Female – With Embeddings
c. Male – Without Embeddings
d. Male – With Embeddings
Second, prediction accuracy is assessed using the class probabilities generated by the transformer-based Zero-Shot classifier. This analysis uses the classes “Personal Growth”, “Community Support”, “Rule Violations”, and “Disruptive Conduct”. Each of these classes is chosen to be indicative of behaviors that are either meaningfully detrimental to community peace and cohesion or supportive of personal growth and community well-being. We use the class probabilities as predictors instead of the above text embeddings. Considering there are only four classes in our specification, there is no need to use LASSO. Instead, we use a logistic regression model. However, since the class probabilities are compositional data, we transform the probabilities into additive log ratios, using disruptive conduct as the reference category. We again obtain substantial improvement in out-of-sample prediction accuracy (29.7% for females and 22% for males) when including class probabilities along with the matrix of observed covariates (Figure 6).
@@
a. Female – Without Class Prob.
b. Female – With Class Prob.
c. Male – Without Class Prob.
d. Male – With Class Prob.
Finally, we provide prediction results on recidivism aggregated by receiver profiles with and without the class probabilities (Figure B7). We find the prediction accuracy improves to the tune of 35% and 23% for the female and male units, respectively.

### v2

@@ 4.6 Network models
In terms of latent variable network models, we first consider the Latent Space Model (LSM), which is a general random graph model that includes both additive and multiplicative latent variables and can also accommodate covariates (Hoff et al.,, 2002; Hoff,, 2021; Ma et al.,, 2020; Li et al.,, 2023). The model is parameterized by $d$-dimensional unknown vectors $\bm{q}_{i}$ and unknown scalars $v_{i}$ for all nodes $i=1,\ldots,n$.
We also assume that we have edge level covariates $\bm{X}=(x_{ij})$ available to us. A network adjacency matrix $\bm{A}=(a_{ij})$ from this model is generated as follows (Li et al.,, 2023; Ma et al.,, 2020):
@@
where $f$ is a family of distributions satisfying certain smoothness conditions, and $\sigma:\mathbb{R}\to\mathbb{R}$ is a known link function. In this paper, we consider a particular sparse version of the model due to Li et al., (2023) that assumes $f$ to be a Bernoulli distribution, link function $\sigma$ to be the logistic function, and there exists a sparsity parameter $\rho_{N}$ such that $\theta_{ij}=logistic({\bm{q}_{i}}^{\prime}\bm{q}_{j}+v_{i}+v_{j}+x_{ij}\beta+\rho_{N})$. Defining $\omega_{N}=\exp(\rho_{N})$, we assume $\omega_{N}\to 0$ and $\omega_{N}=\omega(N^{-1/2})$. We let $\bm{u}_{i}\triangleq(\bm{q}_{i}^{\prime},v_{i})^{\prime}$ denote the $d+1$ dimensional parameter vector containing all latent variables associated with node $i$. A special case of this model is Random Dot Product Graph (RDPG) model, which can be expressed as follows,
@@
where $\rho_{N}$ again controls the sparsity of the model. We will assume $\rho_{N}=\omega(\frac{\log^{4}N}{N})$. Note the expected density (i.e., probability of an edge) scales with $N$ as $\rho_{N}$, or in other words, the expected degree scales as $N\rho_{N}$ in this model, making the model suitable for sparse networks. This model does not account for the additive latent variables $v_{i}$’s but contains $d$-dimensional multiplicative latent variables, and the link function $\sigma(\cdot)$ is specified as identity. Clearly, both these models allow for multidimensional latent variables in the network formation model and are more general than the single latent variable model considered in Johnsson and Moon, (2021). Both models are also capable of modeling sparse networks, which the models in Johnsson and Moon, (2021); Graham, (2017); Auerbach, (2022) are not capable of. We compare the model fit of the latent-space network models with the network model considered in Johnsson and Moon, (2021); Graham, (2017) for the network data in our real-data analysis application (see section 6.2).
We employ the maximum likelihood estimator with Lagrange adjustment in Li et al., (2023) for the estimation of the LSM model. Let $\hat{\bm{U}}$ denote the maximum likelihood estimator over the constrained parameter space
@@
Under the set of assumptions laid out in Li et al., (2023), an upper bound on the error rate of estimating the latent positions in Frobenius norm (Li et al.,, 2023) is $\frac{1}{N}\|\hat{\bm{U}}-\bm{U}\|_{F}^{2}=O_{p}(\frac{1}{N\omega_{N}})$.
Although the LSM described above contains the RDPG model as its special case, there exists a rich amount of results with better convergence rates of the estimate of $\bm{U}$ when one uses spectral embedding to estimate $\bm{U}$ in the case of RDPG model (Athreya et al.,, 2018; Cape et al.,, 2019; Xie and Xu,, 2023; Chang and Paul,, 2024). Specifically, the spectral embedding of the adjacency matrix $A$ can be written as $\hat{\bm{U}}=\bm{U}_{A}|\bm{S}_{A}|^{1/2}$, where $\bm{U}_{A}$ denotes the matrix containing the leading $d$ eigenvectors of $\bm{A}$, which are assumed, without loss of generality, to be ordered by decreasing absolute eigenvalue magnitude. The matrix $|\bm{S}_{A}|$ is a diagonal matrix containing the absolute values of the corresponding eigenvalues. Then, a faster convergence rate than LSM can be obtained in terms of the $2\to\infty$ norm (maximum of row-wise $\ell_{2}$ norms) as $\|\hat{\bm{U}}-\bm{U}\bm{H}\|_{2,\infty}=O_{hp}\left(\frac{\log^{c}N}{N^{1/2}}\right)$ (Rubin-Delanchy et al.,, 2022; Cape et al.,, 2019). Here we mean $X_{n}=O_{hp}(1)$ by: for every $c>0$ there exist constants $M(c),n_{0}(c)>0$ such that $\Pr(|X_{n}|>M)<n^{-c}$ for all $n>n_{0}$.
@@ 4.7 Theoretical results
For a matrix $A$ we denote its spectral norm as $\|A\|=\max_{\|x\|=1}|Ax|$, Frobenius norm as $\|A\|_{F}=\sqrt{\sum_{ij}a_{ij}^{2}}$ and the $2\to\infty$ norm as $\|A\|_{2,\infty}=\max_{i}\sqrt{\sum_{j}a_{ij}^{2}}$. We use the stochastic order notation $o_{p}(1)$ to mean that if $x_{N}=o_{p}(1)$, then $x_{N}\overset{p}{\to}0$. We start with a list of assumptions. The first assumption is a collection of conditions described above in the description of the model, making the model well-behaved and identification possible, similar to those in (Johnsson and Moon,, 2021).
@@
We note again that the network $\bm{A}$ is endogenous by the dependence between $\bm{U}_{i}$ and $\bm{E}_{i}$. A consequence of Assumption 1 is that
@@
The above assumption is an extension of the assumption in Johnsson and Moon, (2021) to multivariate basis functions which says the true component-wise functions $h_{j}(u)$ are well approximated by the sieve approximations for all vectors $u$ and all components $j$.
@@
Assume the assumptions in the section 4.7. Let the graph $\bm{A}$ be generated via either RDPG model with sparsity parameter $\rho_{N}=\omega(\frac{\log^{4}N}{N})$ or LSM with sparsity parameter $\omega_{N}=\omega(N^{-1/2})$ as described in section 4.6, respectively. Further assume that
@@
The proofs of all results are in the Appendix A1. Several additional technical lemmas are needed, which are all stated and proved in the Appendix A1.
Note that the conditions on the graph density indicate that consistent and asymptotically normal estimation is possible even for sparse graphs. Under a binary graph ($a_{ij}$s are either 1 or 0), for example, the first condition in (6) only requires $\delta_{\max}/\delta_{\min}^{2}=o_{p}(1/\sqrt{N})$. A sufficient condition for the second part in (7) can be (since for a binary graph $\|\bm{a}_{i_{1}}\|^{2}=\delta_{i_{1}}$ for any $i_{1}$)
@@
These conditions hold under a moderate expected density of the graph. For example, if we let $\delta_{\min}\asymp\delta_{\max}$ and the average expected density of the graph grows as $O(\frac{1}{\sqrt{N}})$, then all conditions are satisfied. This is in contrast to the results in Johnsson and Moon, (2021), which only hold if the average expected density of the graph grows as $O(1)$, making the graph unrealistically dense.
@@
Let Assumption 3 hold. Under the conditions on RDPG model in Rubin-Delanchy et al., (2022) and sparsity parameter $\rho_{N}=\omega(\frac{\log^{4}N}{N})$, we have
@@
and under the conditions on the sparse LSM model with sparsity parameter $\omega_{N}=\omega(N^{-1/2})$ in Li et al., (2023), we have
@@ 5 Simulations
Table A2.2 documents the performance of the estimator when the network is generated from the RDPG model and the spectral embedding is used to estimate the latent variables. The dimension of latent variables in the outcome model is fixed at 2. There are two correlated outcomes where the data generation follows Equation 1 with endogeneity in network formation. The peer influence parameter matrix is set at $\begin{bmatrix}0.3&0.2\\ 0.25&0.6\end{bmatrix}$ where the diagonal elements capture the direct peer influence spillovers and the off-diagonal elements capture the indirect spillovers in the outcome. The dimension of the coefficients on $h^{Y}(U)$ varies across the three scenarios. For each scenario, the data is generated 100 times, followed by estimation of $\beta_{v}$ using the steps in algorithm 1. The table shows that the mean squared error steadily declines with the number of nodes in the network. This supports the fact that the estimator works well in recovering the true parameter of interest as the sample size increases.
Next, we change the data generation and estimation of latent variables, where latent space models that allow for both multiplicative and additive unobserved variables, along with observed covariates, are used for generating the network. A Projected Gradient Descent (PGD) algorithm with Universal Singular Value Thresholding (USVT) initialization is used to estimate the MLE $\hat{U}$ (Ma et al.,, 2020). Note that even though this solves a non-convex optimization problem, the solution has nice theoretical guarantees which we have used in the proofs above. The covariate that explains the network formation is not part of the data generation for the outcomes. The network formation consists of one additive latent factor, two multiplicative latent factors, and two observed covariates. The outcome data generation consists of $AY,X,AX,h^{Y}(U)]$ where $AY$ is $N\times 2$ matrix, $X$ is $N\times 5$, $AX$ is $N\times 5$ and dimension of $h^{Y}(U)$ varies across the 3 scenarios. The peer influence matrix $D$ used in the data generation for the latent space model with covariates is fixed at $\begin{bmatrix}0.8&0.2\\ 0.3&0.6\end{bmatrix}$. Table A2.3 reports the mean squared error (MSE) across 50 simulations for each scenario as the sample size increases from 100 to 300 to 500. It shows that the MSE consistently declines as the sample size increases for all four parameters of the peer influence $D$ matrix for all three scenarios.
@@ 6.1 Predict Recidivism using Text Messages
a. Female
b. Male
Notes: Five-fold cross-fitting is used to estimate the out-of-sample AUC for all models. Logistic LASSO is used for the prediction models because the text embeddings are high-dimensional.
Figure 5 provides the Receiver Operator Characteristic (ROC) curves and the Area Under the Curve (AUC) for two models, one with both pre-entry covariates ($X$) and embeddings ($T$), and an alternative model without the text embeddings (only $X$) as predictors. Panel (a) shows that the prediction AUC is 30% higher for recidivism in the female unit when we incorporate the LLM-based text embeddings relative to the model with only pre-entry covariates $X$. The dimension reduction is substantial for the former model, as about 26-37 predictors out of text embeddings (768) and pre-entry covariates remain across the five folds in LASSO model fit. Panel (b) compares the prediction accuracy for the male units with and without the embeddings. The prediction accuracy improvement for the males is 30.6%. Figure A2.6 repeats the same analysis but for the case when the text embeddings are aggregated by the receiver ID. We find comparable improvements (32.6% for females and 32% for males) in predictive accuracy when the text embeddings are aggregated by the receiver profiles and included as predictors in the LASSO model along with the pre-entry covariates. Recall from section 3 that the sender profile indicates how an individual interacts with their peers, while the receiver profiles indicate how others perceive an individual.
The out-of-sample AUC values from our model with embeddings are in the range of 0.67-0.70, which is generally considered a good predictive accuracy in the context of recidivism prediction (Laqueur and Copus,, 2024). We further note that our pre-entry covariates include an aggregate measurement known as Level of Service Inventory-Revised (LSI-R), a widely used measure in the context of criminal psychology (Andrews and Bonta,, 2014). The LSI-R survey contains 54 questions pertaining to criminal history, education, employment, family, alcohol and drug problems, living conditions, and emotional health. Several questions that are part of the LSI-R are known to be predictors of recidivism (Dressel and Farid,, 2018). Yet, in our sample, the pre-entry covariates, including LSI-R, were only mildly predictive of recidivism. This could be because LSI-R is an aggregated measure, and we do not observe the component scores in our data. Nevertheless, this underscores the importance of AI-based text measurements as potential predictors of recidivism, especially in low-security correctional facilities, where observed covariates have limited ability to predict recidivism.
a. Female
b. Male
Next, prediction accuracy is assessed using the class probabilities generated by the transformer-based Zero-Shot classifier. This analysis uses the classes “Personal Growth”, “Community Support”, “Rule Violations”, and “Disruptive Conduct”. Each of these classes is chosen to be indicative of behaviors that are either meaningfully detrimental to community peace and cohesion or supportive of personal growth and community well-being. We use the class probabilities as predictors instead of the above text embeddings. Considering there are only four classes in our specification, there is no need to use LASSO. Instead, we use a logistic regression model. However, since the class probabilities are compositional data, we transform the probabilities into additive log ratios, using disruptive conduct as the reference category. We again obtain substantial improvement in out-of-sample prediction accuracy (29.7% for females and 22% for males) when including class probabilities along with the matrix of observed covariates (Figure 6).
@@
Further, in Figure 8(a), we assess the (out-of-sample) predicted probabilities from our compositional logistic regression model with zero-shot classes against the true recidivism outcome. The comparison of the two histograms shows that the logistic model assigns a higher probability to residents who eventually recidivate. In Figure 8(b), we show a calibration plot between the mean predicted probability and the mean observed outcome in 20 quantiles. To do so, we divide the predicted probabilities into 20 quantiles. In each quantile, we compute the mean of the predictions and plot it against the mean of the observed outcomes for individuals in that quantile. If our prediction model is accurate, we expect these points to lie close to a 45-degree line. The figure shows that most points lie close to the red 45-degree line, indicating good calibration of our model’s predictions. Similar exercise is done for the male unit, and the results are provided in Figure A2.5 in the Appendix.
Finally, we provide prediction results on recidivism aggregated by receiver profiles with and without the class probabilities (Figure A2.7). We find that the prediction accuracy improves by 35% and 23% for the female and male units, respectively.
Figure 8: Calibration Plots Predicted Probability and True Recidivism (Female)
a. Histogram True vs. Predicted Outcome
b. Calibration Plot